# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v2

@@ 5 Related Works
With the advent of LLMs such as GPT-4 and Claude, assessing the personality of LLMs has become a popular area of research for the last couple of years (Karra et al., 2022; Jiang et al., 2023b; Miotto et al., 2022; Song et al., 2023; Caron and Srivastava, 2022; Huang et al., 2023; Bodroza et al., 2023; Serapio-García et al., 2023; Pan and Zeng, 2023; Jiang et al., 2024; Noever and Hyams, 2023). Most existing studies typically adopt psychometric questionnaires that are originally proposed for human personality assessment (Pellert et al., 2023, 2022; Serapio-García et al., 2023), such as BFI (John and Srivastava, 1999) or IPIP-NEO (Goldberg et al., 1999), or use machine-generated tests like Anthropic-Eval Perez et al. (2022). However, these tests have self-assessment forms, that lack detailed and varied scenarios when asking about the personality, and are shown to be less reliable due to the sensitivity, occurred by prompt, negation, or order of options (Gupta et al., 2024; Dorner et al., 2023; Frisch and Giulianelli, 2024), resonating our observations. Our TRAIT overcomes the limitations of self-assessment tests, enabling us to measure the personality of LLMs more accurately.
@@ 7 Limitations
In constructing our dataset, we utilize ATOMIC10$\times$ and GPT-4 to generate synthetic data. As is generally known, GPT-4 tends to reflect perspectives more commonly found in the ‘Global North’, and does not represent everyone on Earth equally Manvi et al. (2024). This limitation affects the cultural and social diversity in our dataset and influences the applicability and relevance of our findings to various regions. Additionally, our work focuses only on English language models, presenting a limitation due to our lack of investigation into multilingual models. Multilingual models may behave differently, and understanding these differences could broaden the scope of our findings.
@@ 8 Ethical Considerations
Attributing human-like feelings and mental states to LLMs, a process known as anthropomorphism Airenti (2015), raises ethical concerns about the perception and treatment of these models. While our study aims to assess personality in LLMs, it is crucial to communicate clearly that these models do not possess consciousness or emotions in the human sense. Misinterpreting these traits could lead to unrealistic expectations or ethical dilemmas concerning the rights of AI entities. We advocate for a view of descriptive psychology and try to measure overt patterns in LLM output. Personality should be strictly viewed as a tool for better interaction and alignment with human needs, rather than attributes that confer any form of personhood.
@@ Appendix B List of LLMs Used in Paper
  • Claude-opus Anthropic :
@@
  • Mistral-7B-instruct Jiang et al. (2023a):
@@
  • Llama3-8B-instruct AI@Meta (2024):
@@
  • Llama2-7B-chat Touvron et al. (2023):
@@
  • Gemma-2B-instruct Team et al. (2024):
@@
  • OLMo-7B-instruct Groeneveld et al. (2024):
@@ C.1 Broader Related Works
Collecting high-quality data via manual annotation requires a massive amount of cost and time. For this reason, automatic data generation has been explored, mainly focusing on extractive approaches such as synthetic parsing (Zhang et al., 2021) or pattern matching (Lehmann et al., ; Buck et al., ). More recently, with the emergence of LLMs, utilizing models for data generation (e.g., symbolic knowledge distillation; West et al. (2022)) has been widely explored (Sclar et al., 2022; Bhagavatula et al., 2023; Meng et al., 2022; Liu et al., 2022; Kim et al., 2024). Zheng et al. (2023) and Chen et al. (2022) generate responses with LLM for emotional supportive conversation and task understanding, respectively. Zhou et al. (2022) generate commonsense inferences to improve response generation quality. However, most of the works studying automatic data generation leverage synthesized data for model training. In this work, we collect reliable questionnaires for measuring the personality of LLMs.
@@ C.2 Psychology and Personality
The definition of personality in humans is still controversial Bergner (2020); Mischel et al. (2007); Larsen et al. (2005). We follow Descriptive Psychology, which views the personality as an observable pattern, instead of viewing personality as a causal entity or inner mechanism Bergner (2017); Ossorio (1978, 2006); Schwartz (2019). In other words, just as we perceive someone as assertive who frequently speaks in a commanding tone, descriptive psychology defines personality as observable facts about behaviors. Similarly, we assess the personality of LLMs by analyzing their response patterns given the situations.
@@
With BIG-5 personality dimensions, no single optimal configuration is suggested between various fitness costs and benefits Nettle (2006). The Dark Triad is considered to be lower is better because of socially undesirable qualities Paulhus (2014); Feher and Vernon (2021). For some specific niches in the profession, traits such as (high) Extraversion, Agreeableness, and Openness are sometimes valid predictors of high performance Barrick (2005).
@@ D.1 T-evaluator Training Details
When we train T-evaluator, we built on a Mistral-7B666mistralai/Mistral-7B-v0.1, and use LoRA Hu et al. (2021) for efficient model training. We use lit-gpt AI (2023) framework for model training, using the following hyperparameters: learning rate 3e-4, rank 8, alpha 16, three epochs of training, warmup steps 100, batch size of 256, and do single-gpu training in RTX-3090. We adopt the final checkpoint of iteration.
@@ D.2 Token Probability Measurement
For every question, we adopt a multi-choice QA (MCQA) format with four possible options (i.e., tokens A, B, C, and D followed by the choices), two options labeled with ‘High’ and the other two labeled with ‘Low’. We follow the evaluation procedure of various MCQA benchmarks such as MMLU Hendrycks et al. (2020) which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice; first by assigning ‘A: High, B: Low, C: High, D: Low’ and then reversing the high and low values to ‘A: Low, B: High, C: Low, D: High’. After that, we calculate the average probability of tokens from two arrangements for each option and designate the option with the highest probability as the preferred option by LLM. Finally, the score for each personality trait is evaluated by the ratio of ‘High’ responses to the total number of questions.
@@ E.1.3 Reliability
three different prompt template for each test is presented in Table 25(a) to 27(c).
@@
When we measure semantic similarity, we use BERTScore Zhang et al. (2019) and calculate the retrieval accuracy. Using BERTScore, we retrieve the paraphrased option from the original four options (column ‘Options’). We retrieve paraphrased question from randomly sampled 100 questions that have same personality trait (Column ‘question’). In Figure 9, the accuracy of retrieval task is shown. Our paraphrased sentences show high score of accuracy in the retrieval task, showing that little semantic difference between the original sentence and the paraphrased sentence.
@@ F.1 More LLM Personality Test results on TRAIT
Test Template type Openness Conscientiousness Extraversion Agreeableness Neuroticism Psychopathy Machiavellism Narcissism
GPT-4 Type 1 56.5 93.9 33.7 85.1 23 0.3 11.9 7.7
@@
Type 3 59.9 90.1 38.6 83.7 27 0.5 9.6 8.4
average (std) 58.4 (1.43) 92.6 (1.79) 35.3 (2.36) 85.5 (1.7) 24.4 (1.82) 0.3 (0.16) 11 (1.02) 7.5 (0.78)
Claude-opus Type 1 49.7 91.7 23.65 84.6 25.0 0 7.8 3.8
Type 2 55.1 91.9 24.1 88.3 22.9 0 4.8 1.75
Type 3 58.7 88.7 32.4 87.15 23.2 0 9.3 4.95
average (std) 54.5 (3.7) 90.8 (1.45) 26.7 (4) 86.7 (1.57) 23.7 (0.93) 0 (0) 7.3 (1.89) 3.5 (1.32)
Gemini-1.0-pro Type 1 72.5 95 46.2 87.5 35.3 2.2 33.9 16.4
Type 2 48 84.6 19.6 74.2 20.9 1.1 5.8 4.1
Type 3 60.25 89.8 32.9 80.85 28.1 1.65 19.85 10.25
average (std) 60.3 (10) 89.8 (4.25) 32.9 (10.86) 80.9 (5.43) 28.1 (5.88) 1.7 (0.45) 19.9 (11.47) 10.3 (5.02)
GPT-3.5 Type 1 59 93.8 35.8 75.2 24.2 0.4 17.4 10.9
Type 2 62.7 92.1 30.4 77 25.8 0.2 17.3 8.4
Type 3 67.1 92 46.6 64.1 59.2 28.5 31.3 27.3
average (std) 62.9 (3.31) 92.6 (0.83) 37.6 (6.73) 72.1 (5.7) 36.4 (16.14) 9.7 (13.29) 22 (6.58) 15.5 (8.38)
@@
Type 2 72.2 77.9 58.9 58 19.9 36.5 40 36.9
Type 3 67.4 73.3 50.2 49.9 47.1 51.2 40.3 43
average (std) 69.2 (2.12) 75.6 (1.88) 55.1 (3.65) 53.2 (3.46) 33.9 (11.12) 48.1 (8.49) 42.7 (3.61) 42.1 (3.89)
Llama2-7B-chat Type 1 58 84.2 45.6 73.4 44 23.2 29.9 24
@@
average (std) 74.9 (9.41) 86.2 (4.01) 48.4 (12.02) 71.5 (3.41) 21.7 (1.97) 16.9 (10.77) 20.8 (2.9) 17.5 (6.21)
Llama3-8B-inst Type 1 52.7 88.5 30.3 74.4 30.7 8.6 16.6 9
@@
Type 3 65.4 85.8 43.7 78.8 43.4 19.4 22 15.6
average (std) 57.7 (5.54) 88.6 (2.37) 34.6 (6.46) 76.6 (1.8) 35.8 (5.48) 10.6 (6.52) 18.3 (2.64) 11.8 (2.8)
Tulu2-7B-SFT Type 1 59.9 86 33.4 74.7 18.1 6.8 12.4 8.6
Type 2 62 88.7 33.7 78.1 19.3 4.1 13.3 7.6
@@
average (std) 63.2 (3.34) 85.8 (2.45) 35.3 (2.43) 76 (1.5) 20.2 (2.13) 12.7 (10.31) 14.9 (2.97) 9.8 (2.49)
Tulu2-7B-DPO Type 1 59.8 85.2 35 75.3 20.8 5.4 13 8.8
Type 2 61.4 87.8 33 78.6 20.1 2.7 12 6.9
@@
Type 2 49.4 87.8 32 75.6 33.2 2 13.9 10.2
Type 3 51.8 88.9 31.5 69.9 43.7 15.3 18.1 17
average (std) 49.3 (2.12) 87.8 (0.86) 31.7 (0.22) 72.4 (2.39) 35.6 (5.92) 6.9 (5.95) 15.6 (1.81) 12.7 (3.05)
@@
Type 2 54.7 91.9 30.1 69 42 2.5 17 11
Type 3 59.9 90.2 40.2 66.4 41.4 20.8 20.5 18
average (std) 56.2 (2.6) 90.9 (0.74) 35.2 (4.12) 67.2 (1.25) 40 (2.42) 8.5 (8.7) 18 (1.78) 13.4 (3.23)
@@
average (std) 56.7 (5.44) 60.2 (7.76) 56.1 (3.21) 55.7 (6.91) 40.2 (7.35) 60.1 (5.51) 49.7 (0.33) 54.7 (6.81)
OLMo-7B-instruct Type 1 56 89.1 42.6 67.2 25.9 22.2 16.1 19.1
Type 2 66.3 91.1 39.3 76.2 32 21.3 23.2 15.9
Type 3 64 81.6 51.5 56.7 41.7 74 34.2 35.3
average (std) 62.1 (4.41) 87.3 (4.09) 44.5 (5.15) 66.7 (7.97) 33.2 (6.51) 39.2 (24.63) 24.5 (7.45) 23.4 (8.49)
Gemma-2B Type 1 59 77.6 49.9 52 42.7 39.9 37.3 45.9
Type 2 74.3 81 55.1 74.3 27.7 35.3 29.4 25.4
Type 3 66.2 58 60.1 49.2 17.3 64.1 37.7 50.6
average (std) 66.5 (6.25) 72.2 (10.14) 55 (4.16) 58.5 (11.23) 29.2 (10.43) 46.4 (12.63) 34.8 (3.82) 40.6 (10.94)
@@
average (std) 70.4 (2.61) 89 (6.2) 42.1 (7.21) 70.5 (2.53) 32.3 (2.21) 29.8 (11.26) 18.9 (3.04) 24.8 (6.17)
Qwen 1.5-7B-Chat Type 1 60.1 94.4 33.7 85.7 20.9 0.5 14.8 9
Type 2 60.2 93.9 31.5 86.8 23 1.7 17 8.7
@@
Model Trait Level Template Type
Type 1 Type 2 Type 3 Mean Std
GPT-4 Openness High 90.4 95.7 97.1 94.4 2.89
Low 1.5 0.7 0.6 0.9 0.40
Conscientiousness High 99.0 99.2 99 99.1 0.09
Low 12.8 4.1 1.3 6.1 4.90
Extraversion High 90.3 97.2 99.5 95.7 3.91
Low 4.6 3.0 2.0 3.2 1.07
Agreeableness High 98.0 98.1 97.3 97.8 0.36
Low 0.2 0.0 0.2 0.1 0.09
Neuroticism High 75.0 87.5 94.3 85.6 7.99
Low 4.6 3.0 2.1 3.2 1.03
Psychopathy High 37.3 80.0 99.7 72.3 26.05
Low 0.0 0.0 0.0 0.0 0.00
Machiavellianism High 98.5 99.1 98.7 98.8 0.25
Low 3.1 3.0 2.0 2.7 0.50
Narcissism High 99.1 99.5 99.5 99.4 0.19
Low 2.1 2.1 2.5 2.2 0.19
GPT-3.5 Openness High 92.8 95.7 94.0 94.2 1.19
Low 1.6 21.6 57.1 26.8 22.95
Conscientiousness High 98.4 98.0 98.7 98.4 0.29
Low 5.7 24.7 63.4 31.3 24.01
Extraversion High 85.1 94.6 96.5 92.1 4.99
Low 3.5 13.2 25.2 14.0 8.88
Agreeableness High 91.7 88.9 86.3 89.0 2.21
Low 9.3 5.5 6.7 7.2 1.59
Neuroticism High 78.2 81.9 93.8 84.6 6.66
Low 9.1 23.8 59.7 30.9 21.25
Psychopathy High 97.4 99.5 99.9 98.9 1.10
Low 0.0 0.5 34.5 11.7 16.15
Machiavellianism High 94.9 98.9 98.3 97.4 1.76
Low 2.8 6.6 17.9 9.1 6.41
Narcissism High 90.1 98.9 97.9 95.6 3.93
Low 0.9 1.8 12.6 5.1 5.32
Mistral-7B-instruct Openness High 70.6 78.4 84.5 77.8 5.69
Low 11.5 1.9 6.3 6.6 3.92
Conscientiousness High 93.0 94.3 96.3 94.5 1.36
Low 48.2 13.3 40.3 33.9 14.94
Extraversion High 67.5 76.3 88.3 77.4 8.52
Low 5.3 3.3 1.8 3.5 1.43
Agreeableness High 83.6 89.6 86.7 86.6 2.45
Low 15.5 8.8 9.4 11.2 3.03
Neuroticism High 55.8 60.4 71.1 62.4 6.41
Low 17.4 11.7 14.6 14.6 2.33
Psychopathy High 56.7 90.8 81.0 76.2 14.33
Low 3.3 0.8 2.2 2.1 1.02
Machiavellianism High 74.0 77.9 77.2 76.4 1.70
Low 10.2 6.6 5.6 7.5 1.98
Narcissism High 64.6 78.2 74.3 72.4 5.72
Low 3.7 2.0 3.5 3.1 0.76
Llama2-7B-chat Openness High 87.8 83.2 96.7 89.2 5.60
Low 62.4 44.0 54.4 53.6 7.53
Conscientiousness High 80.1 67.3 96.3 81.2 11.87
Low 64.9 32.2 43.5 46.9 13.56
Extraversion High 81.2 85.7 95.5 87.5 5.97
Low 27.0 37.4 34.6 33.0 4.39
Agreeableness High 76.3 81.5 93.9 83.9 7.38
Low 42.5 32.4 31.0 35.3 5.12
Neuroticism High 53.4 38.2 84.4 58.7 19.23
Low 12.3 10.0 9.7 10.7 1.16
Psychopathy High 56.2 63.3 97.2 72.2 17.89
Low 12.1 14.6 39.4 22.0 12.32
Machiavellianism High 73.3 65.7 92.4 77.1 11.23
Low 20.6 19.0 48.8 29.5 13.69
Narcissism High 64.5 70.2 89.2 74.6 10.56
Low 14.7 13.7 31.0 19.8 7.93
Table 12: Fine-grained results of Figure 5, the prompted models’ personality scores on TRAIT.
In Table 11, we show results from a total 19 models when testing with TRAIT. We report the average scores with three different prompt types and standard deviations. In Table 12, four model results when testing with TRAIT are shown. We also report the average scores with three different prompt types and standard deviations.

### v3

@@ 5 Related Works
With the advent of LLMs such as GPT-4 and Claude, assessing the personality of LLMs has become a popular area of research for the last couple of years (karra2022estimating; jiang2023personallm; miotto2022gpt; song2023have; caron2022identifying; huang2023chatgpt; bodroza2023personality; serapiogarcía2023personality; pan2023llms; jiang2024evaluating; noever2023ai). Most existing studies typically adopt psychometric questionnaires that are originally proposed for human personality assessment (pellert2023ai; pellert2022large; serapiogarcía2023personality), such as BFI (John1999TheBF) or IPIP-NEO (goldberg1999broad), or use machine-generated tests like Anthropic-Eval perez2022discovering. However, these tests have self-assessment forms, that lack detailed and varied scenarios when asking about the personality, and are shown to be less reliable due to the sensitivity, occurred by prompt, negation, or order of options (gupta2024selfassessment; dorner2023personality; frisch2024llm), resonating our observations. Our TRAIT overcomes the limitations of self-assessment tests, enabling us to measure the personality of LLMs more accurately.
@@ 7 Limitations
In constructing our dataset, we utilize ATOMIC10$\times$ and GPT-4 to generate synthetic data. As is generally known, GPT-4 tends to reflect perspectives more commonly found in the ‘Global North’, and does not represent everyone on Earth equally manvi2024large. This limitation affects the cultural and social diversity in our dataset and influences the applicability and relevance of our findings to various regions. Additionally, our work focuses only on English language models, presenting a limitation due to our lack of investigation into multilingual models. Multilingual models may behave differently, and understanding these differences could broaden the scope of our findings.
@@ 8 Ethical Considerations
Attributing human-like feelings and mental states to LLMs, a process known as anthropomorphism airenti2015cognitive, raises ethical concerns about the perception and treatment of these models. While our study aims to assess personality in LLMs, it is crucial to communicate clearly that these models do not possess consciousness or emotions in the human sense. Misinterpreting these traits could lead to unrealistic expectations or ethical dilemmas concerning the rights of AI entities. We advocate for a view of descriptive psychology and try to measure overt patterns in LLM output. Personality should be strictly viewed as a tool for better interaction and alignment with human needs, rather than attributes that confer any form of personhood.
@@ Appendix B List of LLMs Used in Paper
  • Claude-opus claude3:
@@
  • Mistral-7B-instruct jiang2023mistral:
@@
  • Llama3-8B-instruct llama3modelcard:
@@
  • Llama2-7B-chat touvron2023llama:
@@
  • Gemma-2B-instruct gemmateam2024gemma:
@@
  • OLMo-7B-instruct groeneveld2024olmo:
@@ C.1 Broader Related Works
Collecting high-quality data via manual annotation requires a massive amount of cost and time. For this reason, automatic data generation has been explored, mainly focusing on extractive approaches such as synthetic parsing (zhang2021transomcs) or pattern matching (lehmanndbpedia; buckn). More recently, with the emergence of LLMs, utilizing models for data generation (e.g., symbolic knowledge distillation; west2022symbolic) has been widely explored (sclar2022referee; bhagavatula2023i2d2; meng2022generating; liu2022wanli; kim2024pearl). zheng2023augesc and chen2022weakly generate responses with LLM for emotional supportive conversation and task understanding, respectively. zhou2022reflect generate commonsense inferences to improve response generation quality. However, most of the works studying automatic data generation leverage synthesized data for model training. In this work, we collect reliable questionnaires for measuring the personality of LLMs.
@@ C.2 Psychology and Personality
The definition of personality in humans is still controversial bergner2020personality; mischel2007introduction; larsen2005personality. We follow Descriptive Psychology, which views the personality as an observable pattern, instead of viewing personality as a causal entity or inner mechanism bergner2017person; ossorio1978personality; ossorio2006behavior; schwartz2019descriptive. In other words, just as we perceive someone as assertive who frequently speaks in a commanding tone, descriptive psychology defines personality as observable facts about behaviors. Similarly, we assess the personality of LLMs by analyzing their response patterns given the situations.
@@
With BIG-5 personality dimensions, no single optimal configuration is suggested between various fitness costs and benefits nettle2006evolution. The Dark Triad is considered to be lower is better because of socially undesirable qualities paulhus2014toward; feher2021looking. For some specific niches in the profession, traits such as (high) Extraversion, Agreeableness, and Openness are sometimes valid predictors of high performance barrick2005yes.
@@ D.1 T-evaluator Training Details
When we train T-evaluator, we built on a Mistral-7B666mistralai/Mistral-7B-v0.1, and use LoRA hu2021lora for efficient model training. We use lit-gpt litgpt-2023 framework for model training, using the following hyperparameters: learning rate 3e-4, rank 8, alpha 16, three epochs of training, warmup steps 100, batch size of 256, and do single-gpu training in RTX-3090. We adopt the final checkpoint of iteration.
@@ D.2 Token Probability Measurement
For every question, we adopt a multi-choice QA (MCQA) format with four possible options (i.e., tokens A, B, C, and D followed by the choices), two options labeled with ‘High’ and the other two labeled with ‘Low’. We follow the evaluation procedure of various MCQA benchmarks such as MMLU hendrycks2020measuring which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice; first by assigning ‘A: High, B: Low, C: High, D: Low’ and then reversing the high and low values to ‘A: Low, B: High, C: Low, D: High’. After that, we calculate the average probability of tokens from two arrangements for each option and designate the option with the highest probability as the preferred option by LLM. Finally, the score for each personality trait is evaluated by the ratio of ‘High’ responses to the total number of questions.
@@ E.1.3 Reliability
three different prompt template for each test is presented in Table 17(a) to 19(c).
@@
When we measure semantic similarity, we use BERTScore zhang2019bertscore and calculate the retrieval accuracy. Using BERTScore, we retrieve the paraphrased option from the original four options (column ‘Options’). We retrieve paraphrased question from randomly sampled 100 questions that have same personality trait (Column ‘question’). In Figure 9, the accuracy of retrieval task is shown. Our paraphrased sentences show high score of accuracy in the retrieval task, showing that little semantic difference between the original sentence and the paraphrased sentence.
@@ F.1 More LLM Personality Test results on TRAIT
Test Template Ope. Con. Ext. Agr. Neu. Psy. Mac. Nar.
GPT-4 Type 1 56.5 93.9 33.7 85.1 23.0 0.3 11.9 7.7
@@
Type 3 59.9 90.1 38.6 83.7 27.0 0.5 9.6 8.4
Mean 58.4 92.6 35.3 85.5 24.4 0.3 11.0 7.5
Std 1.43 1.79 2.36 1.70 1.82 0.16 1.02 0.78
Claude-opus Type 1 49.7 91.7 23.7 84.6 25.0 0.0 7.8 3.8
Type 2 55.1 91.9 24.1 88.3 22.9 0.0 4.8 1.8
Type 3 58.7 88.7 32.4 87.2 23.2 0.0 9.3 5.0
Mean 54.5 90.8 26.7 86.7 23.7 0.0 7.3 3.5
Std 3.70 1.45 4.00 1.57 0.93 0.00 1.89 1.32
Gemini-1.0-pro Type 1 72.5 95.0 46.2 87.5 35.3 2.2 33.9 16.4
Type 2 48.0 84.6 19.6 74.2 20.9 1.1 5.8 4.1
Type 3 60.3 89.8 32.9 80.9 28.1 1.7 19.9 10.3
Mean 60.3 89.8 32.9 80.9 28.1 1.7 19.9 10.3
Std 10.00 4.25 10.86 5.43 5.88 0.45 11.47 5.02
GPT-3.5 Type 1 59.0 93.8 35.8 75.2 24.2 0.4 17.4 10.9
Type 2 62.7 92.1 30.4 77.0 25.8 0.2 17.3 8.4
Type 3 67.1 92.0 46.6 64.1 59.2 28.5 31.3 27.3
Mean 62.9 92.6 37.6 72.1 36.4 9.7 22.0 15.5
Std 3.31 0.83 6.73 5.70 16.14 13.29 6.58 8.38
@@
Type 2 72.2 77.9 58.9 58.0 19.9 36.5 40.0 36.9
Type 3 67.4 73.3 50.2 49.9 47.1 51.2 40.3 43.0
Mean 69.2 75.6 55.1 53.2 33.9 48.1 42.7 42.1
Std 2.12 1.88 3.65 3.46 11.12 8.49 3.61 3.89
Llama2-7B-chat Type 1 58.0 84.2 45.6 73.4 44.0 23.2 29.9 24.0
@@
Mean 74.9 86.2 48.4 71.5 21.7 16.9 20.8 17.5
Std 9.41 4.01 12.02 3.41 1.97 10.77 2.90 6.21
Llama3-8B-inst Type 1 52.7 88.5 30.3 74.4 30.7 8.6 16.6 9.0
@@
Type 3 65.4 85.8 43.7 78.8 43.4 19.4 22.0 15.6
Mean 57.7 88.6 34.6 76.6 35.8 10.6 18.3 11.8
Std 5.54 2.37 6.46 1.80 5.48 6.52 2.64 2.80
Tulu2-7B-SFT Type 1 59.9 86.0 33.4 74.7 18.1 6.8 12.4 8.6
Type 2 62.0 88.7 33.7 78.1 19.3 4.1 13.3 7.6
@@
Mean 63.2 85.8 35.3 76.0 20.2 12.7 14.9 9.8
Std 3.34 2.45 2.43 1.50 2.13 10.31 2.97 2.49
Table 11: Fine-grained personality scores of various models on TRAIT.
Test Template Ope. Con. Ext. Agr. Neu. Psy. Mac. Nar.
Tulu2-7B-DPO Type 1 59.8 85.2 35.0 75.3 20.8 5.4 13 8.8
Type 2 61.4 87.8 33.0 78.6 20.1 2.7 12 6.9
@@
Type 2 49.4 87.8 32.0 75.6 33.2 2.0 13.9 10.2
Type 3 51.8 88.9 31.5 69.9 43.7 15.3 18.1 17.0
Mean 49.3 87.8 31.7 72.4 35.6 6.9 15.6 12.7
Std 2.12 0.86 0.22 2.39 5.92 5.95 1.81 3.05
@@
Type 2 54.7 91.9 30.1 69.0 42.0 2.5 17.0 11.0
Type 3 59.9 90.2 40.2 66.4 41.4 20.8 20.5 18.0
Mean 56.2 90.9 35.2 67.2 40.0 8.5 18.0 13.4
Std 2.60 0.74 4.12 1.25 2.42 8.70 1.78 3.23
@@
Mean 56.7 60.2 56.1 55.7 40.2 60.1 49.7 54.7
Std 5.44 7.76 3.21 6.91 7.35 5.51 0.33 6.81
OLMo-7B-instruct Type 1 56.0 89.1 42.6 67.2 25.9 22.2 16.1 19.1
Type 2 66.3 91.1 39.3 76.2 32.0 21.3 23.2 15.9
Type 3 64.0 81.6 51.5 56.7 41.7 74.0 34.2 35.3
Mean 62.1 87.3 44.5 66.7 33.2 39.2 24.5 23.4
Std 4.41 4.09 5.15 7.97 6.51 24.63 7.45 8.49
Gemma-2B Type 1 59.0 77.6 49.9 52.0 42.7 39.9 37.3 45.9
Type 2 74.3 81.0 55.1 74.3 27.7 35.3 29.4 25.4
Type 3 66.2 58.0 60.1 49.2 17.3 64.1 37.7 50.6
Mean 66.5 72.2 55.0 58.5 29.2 46.4 34.8 40.6
Std 6.25 10.14 4.16 11.23 10.43 12.63 3.82 10.94
@@
Mean 70.4 89.0 42.1 70.5 32.3 29.8 18.9 24.8
Std 2.61 6.20 7.21 2.53 2.21 11.26 3.04 6.17
Qwen 1.5-7B-Chat Type 1 60.1 94.4 33.7 85.7 20.9 0.5 14.8 9.0
Type 2 60.2 93.9 31.5 86.8 23.0 1.7 17.0 8.7
@@
Model Trait (High / Low) Type1 Type2 Type3 Mean Std
GPT-4 Ope. (High / Low) 90.4 / 1.5 95.7 / 0.7 97.1 / 0.6 94.4 / 0.9 2.89 / 0.40
Con. (High / Low) 99.0 / 12.8 99.2 / 4.1 99.0 / 1.3 99.1 / 6.1 0.09 / 4.90
Ext. (High / Low) 90.3 / 4.6 97.2 / 3.0 99.5 / 2.0 95.7 / 3.2 3.91 / 1.07
Agr. (High / Low) 98.0 / 0.2 98.1 / 0.0 97.3 / 0.2 97.8 / 0.1 0.36 / 0.09
Neu. (High / Low) 75.0 / 4.6 87.5 / 3.0 94.3 / 2.1 85.6 / 3.2 7.99 / 1.03
Psy. (High / Low) 37.3 / 0.0 80.0 / 0.0 99.7 / 0.0 72.3 / 0.0 26.05 / 0.00
Mac. (High / Low) 98.5 / 3.1 99.1 / 3.0 98.7 / 2.0 98.8 / 2.7 0.25 / 0.50
Nar. (High / Low) 99.1 / 2.1 99.5 / 2.1 99.5 / 2.5 99.4 / 2.2 0.19 / 0.19
GPT-3.5 Ope. (High / Low) 92.8 / 1.6 95.7 / 21.6 94.0 / 57.1 94.2 / 26.8 1.19 / 22.95
Con. (High / Low) 98.4 / 5.7 98.0 / 24.7 98.7 / 63.4 98.4 / 31.3 0.29 / 24.01
Ext. (High / Low) 85.1 / 3.5 94.6 / 13.2 96.5 / 25.2 92.1 / 14.0 4.99 / 8.88
Agr. (High / Low) 91.7 / 9.3 88.9 / 5.5 86.3 / 6.7 89.0 / 7.2 2.21 / 1.59
Neu. (High / Low) 78.2 / 9.1 81.9 / 23.8 93.8 / 59.7 84.6 / 30.9 6.66 / 21.25
Psy. (High / Low) 97.4 / 0.0 99.5 / 0.5 99.9 / 34.5 98.9 / 11.7 1.10 / 16.15
Mac. (High / Low) 94.9 / 2.8 98.9 / 6.6 98.3 / 17.9 97.4 / 9.1 1.76 / 6.41
Nar. (High / Low) 90.1 / 0.9 98.9 / 1.8 97.9 / 12.6 95.6 / 5.1 3.93 / 5.32
Mistral-7B-instruct Ope. (High / Low) 70.6 / 11.5 78.4 / 1.9 84.5 / 6.3 77.8 / 6.6 5.69 / 3.92
Con. (High / Low) 93.0 / 48.2 94.3 / 13.3 96.3 / 40.3 94.5 / 33.9 1.36 / 14.94
Ext. (High / Low) 67.5 / 5.3 76.3 / 3.3 88.3 / 1.8 77.4 / 3.5 8.52 / 1.43
Agr. (High / Low) 83.6 / 15.5 89.6 / 8.8 86.7 / 9.4 86.6 / 11.2 2.45 / 3.03
Neu. (High / Low) 55.8 / 17.4 60.4 / 11.7 71.1 / 14.6 62.4 / 14.6 6.41 / 2.33
Psy. (High / Low) 56.7 / 3.3 90.8 / 0.8 81.0 / 2.2 76.2 / 2.1 14.33 / 1.02
Mac. (High / Low) 74.0 / 10.2 77.9 / 6.6 77.2 / 5.6 76.4 / 7.5 1.70 / 1.98
Nar. (High / Low) 64.6 / 3.7 78.2 / 2.0 74.3 / 3.5 72.4 / 3.1 5.72 / 0.76
Llama2-7B-chat Ope. (High / Low) 87.8 / 62.4 83.2 / 44.0 96.7 / 54.4 89.2 / 53.6 5.60 / 7.53
Con. (High / Low) 80.1 / 64.9 67.3 / 32.2 96.3 / 43.5 81.2 / 46.9 11.87 / 13.56
Ext. (High / Low) 81.2 / 27.0 85.7 / 37.4 95.5 / 34.6 87.5 / 33.0 5.97 / 4.39
Agr. (High / Low) 76.3 / 42.5 81.5 / 32.4 93.9 / 31.0 83.9 / 35.3 7.38 / 5.12
Neu. (High / Low) 53.4 / 12.3 38.2 / 10.0 84.4 / 9.7 58.7 / 10.7 19.23 / 1.16
Psy. (High / Low) 56.2 / 12.1 63.3 / 14.6 97.2 / 39.4 72.2 / 22.0 17.89 / 12.32
Mac. (High / Low) 73.3 / 20.6 65.7 / 19.0 92.4 / 48.8 77.1 / 29.5 11.23 / 13.69
Nar. (High / Low) 64.5 / 14.7 70.2 / 13.7 89.2 / 31.0 74.6 / 19.8 10.56 / 7.93
Table 13: Prompted models’ fine-grained personality scores on TRAIT, with High/Low merged into one row.
In Table 11 and 12, we show results from a total 19 models when testing with TRAIT. We report the average scores with three different prompt types and standard deviations. In Table 13, four model results when testing with TRAIT are shown. We also report the average scores with three different prompt types and standard deviations.